Given the area in the first quadrant bounded by x² = = 8y, the line


y-2 and the y-axis. What is the volume generated when this area is


revolved about the line y - 2 = 0?

1
Expert's answer
2022-06-07T12:09:25-0400
"y=2=>x^2=8(2)=>x=\pm4"

"V=\pi \displaystyle\int_{-4}^{4}(2-\dfrac{x^2}{8})^2dx"


"V=\pi \displaystyle\int_{-4}^{4}(4-\dfrac{x^2}{2}+\dfrac{x^4}{64}))dx"




"=\pi[4x-\dfrac{x^3}{6}+\dfrac{x^5}{320}]\begin{matrix}
 4\\
 -4
\end{matrix}"




"=\pi(16-\dfrac{64}{6}+\dfrac{1024}{320}-(-16+\dfrac{64}{6}-\dfrac{1024}{320}))"




"=\dfrac{256\pi}{15}({units}^3)"